⋃ 𝐴𝑖


∝


1 and β‹‚ 𝐴𝑖


∞


1


a) 𝐴𝑖 = {𝑖, 𝑖 + 1,𝑖 + 2, … }


b) 𝐴𝑖 = {0,𝑖}


c) 𝐴𝑖 = {βˆ’π‘–, βˆ’ 𝑖 + 1, … , βˆ’1, 0, 1, … , 𝑖 βˆ’ 1, 𝑖}


d) 𝐴𝑖 = {βˆ’π‘–, 𝑖}

1
Expert's answer
2021-11-24T10:36:38-0500

Let us find "\bigcup\limits_{i=1}^{\infty} A_i" and "\bigcap\limits_{i=1}^{\infty} A_i" for the following sets "A_i."


a) If "𝐴_𝑖 = \{𝑖, 𝑖 + 1,𝑖 + 2, … \}" then "\bigcup\limits_{i=1}^{\infty} A_i=A_1=\mathbb N" and "\bigcap\limits_{i=1}^{\infty} A_i=\emptyset."


b) For "𝐴_𝑖 = \{0,𝑖\}" we get that "\bigcup\limits_{i=1}^{\infty} A_i=\{0,1,2,...\}=\mathbb N_0" and "\bigcap\limits_{i=1}^{\infty} A_i=\{0\}."


c) If "𝐴_𝑖 = \{−𝑖, − 𝑖 + 1, … , −1, 0, 1, … , 𝑖 − 1, 𝑖\}" then "\bigcup\limits_{i=1}^{\infty} A_i=\mathbb Z" and "\bigcap\limits_{i=1}^{\infty} A_i=A_1=\{-1,0,1\}."


d) For "𝐴_𝑖 = \{−𝑖, 𝑖\}" we get that "\bigcup\limits_{i=1}^{\infty} A_i=\mathbb Z\setminus\{0\}" and "\bigcap\limits_{i=1}^{\infty} A_i=\emptyset."